# Mitigating Structural Noise in Low-Resource S2TT: An Optimized Cascaded Nepali-English Pipeline with Punctuation Restoration

### v1

@@
Abstract
This paper presents and evaluates an optimized cascaded Nepali speech-to-English text translation (S2TT) system, focusing on mitigating structural noise introduced by Automatic Speech Recognition (ASR). We first establish highly proficient ASR and NMT components: a Wav2Vec2-XLS-R-300m model achieved a state-of-the-art 2.72% CER on OpenSLR-54, and a multi-stage fine-tuned MarianMT model reached a 28.32 BLEU score on the FLORES-200 benchmark. We empirically investigate the influence of punctuation loss, demonstrating that unpunctuated ASR output significantly degrades translation quality, causing a massive 20.7% relative BLEU drop on the FLORES benchmark. To overcome this, we propose and evaluate an intermediate Punctuation Restoration Module (PRM). The final S2TT pipeline was tested across three configurations on a custom dataset. The optimal configuration, which applied the PRM directly to ASR output, achieved a 4.90 BLEU point gain over the direct ASR-to-NMT baseline (BLEU 36.38 vs. 31.48). This improvement was validated by human assessment, which confirmed the optimized pipeline’s superior Adequacy (3.673) and Fluency (3.804). This work validates that targeted punctuation restoration is the most effective intervention for mitigating structural noise in the Nepali S2TT pipeline. It establishes an optimized baseline and demonstrates a critical architectural insight for developing cascaded speech translation systems for similar low-resource languages.
Mitigating Structural Noise in Low-Resource S2TT: An Optimized Cascaded Nepali-English Pipeline with Punctuation Restoration
Tangsang Chongbang  Pranesh Pyara Shrestha   Amrit Sarki   Anku Jaiswal Department of Electronics and Computer Engineering, Pulchowk Campus Institute of Engineering, Tribhuvan University, Nepal {077bei047.tangsang, 077bei030.pranesh, 077bei049.amrit, anku.jaiswal}@pcampus.edu.np
Keywords: Speech-to-Text Translation, Low-Resource Languages, Nepali, ASR, NMT, Nepali-to-English, Punctuation Restoration, Wav2Vec2, MarianMT, Cascaded Pipeline.
1 Introduction
@@
Nepali, spoken by approximately 19 million native speakers, and an additional 14 million second-language speakers Eberhard et al. (2024), presents a range of challenges for speech translation. These include limited annotated data, frequent code-switching with English Gurung (2019), regional dialect variation, and sociolinguistic features such as complex honorifics without direct English equivalents (e.g., t vs. tpA` both translate as “you” but differ in levels of formality). These characteristics complicate both ASR and MT for Nepali.
End-to-end direct Speech-to-Speech translation (S2ST) models are gaining attention in recent years, for their advantages of architectural simplicity, and potential to reduce information loss, and error propagation Bentivogli et al. (2021). However, such approaches demand large-scale, parallel speech and text data Sarim et al. (2025), which makes it impractical for Nepali and most LRLs. In contrast, cascaded pipelines, which first transcribe speech with ASR and then translate it with MT, offer a more practical alternative in low-resource settings. Their modularity allows optimization of each component, resulting in more powerful performance Sarim et al. (2025). Modern self-supervised ASR models like Wav2Vec2 have shown strong performance even with limited labeled data Yi et al. (2020), and multilingual NMT models like MarianMT and NLLB can be adapted even with modest amounts of LRL data to achieve good performance Liu et al. (2020); Verma et al. (2022). Consequently, for low-resource scenarios, cascaded pipelines offer greater data efficiency.
However, cascaded systems introduce their own unique challenges. Error propagation is the most prominent example. A particularly underexplored issue in $\text{Nepali}\leftrightarrow\text{English}$ literature is the lack of punctuation in ASR outputs. While ASR models primarily focus on word accuracy, NMT models are trained on punctuated text, relying on punctuation marks for cues such as sentence boundaries, clause separation, emphasis, disambiguation. Their presence or absence can have significant impact on translation quality Jwalapuram (2023).
@@ 2 Related Work
2.1 Automatic Speech Recognition (ASR) for Low-Resource Languages
@@
Traditional ASR approaches for Nepali relied on Hidden Markov Models (HMMs), as in the early work by Ssarma et al. (2017), which achieved 74.49% accuracy for single-word inputs and 55.55% for three-word phrase inputs Nepali recognition using Voice Activity Detection (VAD), primarily limited by their ability to model complex acoustic variations effectively. The advent of deep learning introduced more sophisticated models like RNN-CTC and CNN-RNN hybrids. Regmi et al. (2019) implemented an RNN-CTC model trained on a 1,320-word custom dataset, that achieved a 34% CER, but struggled with speaker variability and generalization. Bhatta et al. (2020) reported 1.83% CER and 11% WER using a CNN-GRU architecture on the OpenSLR43 text-to-speech (TTS) dataset. However, the studio-quality nature of OpenSLR43 limits real-world applicability. Other efforts, such as Dhakal et al. (2022), utilized bidirectional LSTM paired with ResNet and one-dimensional CNN to achieve a 17.06% CER on OpenSLR dataset. While these studies laid foundational groundwork for Nepali ASR, the models were less capable of generalizing to natural speech variations.
Recent self-supervised models like Wav2Vec2 and Whisper have reshaped ASR for LRLs. By leveraging large-scale unlabeled data for pre-training, they enable better generalization with limited examples Zhu et al. (2021); Hsu et al. (2024); Fatehi (2023). Our work builds upon this paradigm by utilizing the Wav2Vec2-XLS-R-300m model on OpenSLR54 and Common Voice v17. This approach enables us to establish a competitive performance benchmark for Nepali ASR.
2.2 Machine Translation (MT) for Low-Resource Languages
Similar to ASR, NMT for LRLs is constrained by lack of high-quality parallel corpora and linguistic complexities. Traditional Statistical Machine Translation (SMT) and rule-based methods dominated early efforts, but transformer-based NMT has since become the standard. However, NMT models require large datasets and are sensitive to rare words, long sentences, domain mismatch and word-alignment issues Koehn and Knowles (2017).
To address these limitations, researchers have explored transfer learning, backtranslation, data augmentation, and pivot languages Haddow et al. (2022); Talwar and Laasri (2025); Zoph et al. (2016). For instance, MarianMT and NLLB are pretrained multilingual models that support rapid adaptation to LRLs, even with modest amounts of data Liu et al. (2020); Verma et al. (2022). Recent work by Verma et al. (2022) demonstrated the effectiveness of multi-stage fine-tuning strategy involving multilingual pre-training followed by language-pair specific fine-tuning, followed by domain fine-tuning.
@@ 2.2 Machine Translation (MT) for Low-Resource Languages
2.3 Challenges in ASR-MT Integration and Punctuation Restoration
Cascaded ASR-MT pipelines, while practical for LRLs, are inherently susceptible to error propagation, where transcription errors negatively affect downstream translation. Since ASR models like Wav2Vec2 primarily focus on word accuracy, and are not inherently trained to predict punctuation, the resultant degradation from directly feeding the unpunctuated ASR output to NMT models is an under-explored issue, especially for LRLs. Punctuation conveys structural and semantic cues like sentence boundaries, emphasis, disambiguation, which NMT models heavily rely on for accurate translation. The work by Jwalapuram (2023) quantified the effects of punctuation on MT specifically for German-English, Japanese-English and Ukrainian-English language pairs and concluded that models are heavily sensitive to punctuation.
@@ 3 Methodology
3.1 Overall System Architecture
@@
Figure 1: Cascaded Nepali-to-English Speech-To-Text-Translation System Architecture
3.2 Datasets
@@ 3.2.1 ASR Datasets
Table 1: Nepali ASR datasets used for fine-tuning.
Dataset Type Total Utterances/Clips Total Hours (Validated) Used Utterances
Common Voice v17a Crowdsourced 742 (validated) + 595 (internally reviewed ‘other’ split) = 1,337 1 hour (validated) 1,337
OpenSLR54b Scripted/ASR Training $\sim$157,000 (initial) → 148,188 (post-numerals-filtering) $\sim$143.6 hours (pre-5s filter) 136,095 (post-5s filter)
From the Common Voice v17 (ne-NP) dataset, we used 1,337 utterances (approximately 2 hours) from 32 speakers. We combined the validated split with internally reviewed and validated samples from the other split for initial fine-tuning. All audio files were uniformly resampled to 16 kHz.
For OpenSLR54, utterances containing numerals were removed to ensure the model outputs numbers in words rather than digits, since inconsistent transcriptions were common. For example, the numeral 007 appeared as "Ejro Ejro sAt" in one recording, "sΔ Σy sΔ Σy sAt" in another, and simply "sAt" elsewhere. To avoid confusing the model, all such utterances were excluded.
@@
3.2.2 NMT Datasets and Multi-Stage Strategy
@@
Table 2: Corpora used at each stage of NMT training and their filtering/rationale.
Stage Corpus Type Size (Sentence Pairs) Primary Source / Filtering Rationale
Stage 1: Foundational Filtered Web crawled NLLB Corpusa 708,000 Initial 19.6M web-crawled NLLB corpus pairs filtered by LabSE cosine similarity (>>0.80), and removal of Nepali numerals (to maintain consistency). Knowledge Transfer: To learn fundamental translation nuances from filtered, noisy web data.
Stage 2: Expansion Synthetic Corpora 5.02 million Combination of 1.6M synthetic pairsDuwal and Bal (2019)b and 3.42M pre-training pairsDuwal et al. (2024)c filtered using a chrF++ cut-off of 50 (translated with 8-bit NLLB/IndicTrans2). Coverage Expansion: Increase exposure to diverse structures and vocabularies using high-volume, quality-filtered synthetic data.
Stage 3: Refinement Manually translated and validated high-Quality corporad,e 210,875 OPUS (GNOME/KDE/Ubuntu), Bible Translation, Global Voices Parallel Corpus, Penn Treebank + Acharya and Bal (2018), Nepal Law Commission, Easy Bible, Nepal Budget Speech 2081/82, Shirish ko Fool Translation. Quality Refinement: Final tuning for translation fluency and accuracy on high-fidelity human-translated pairs.
3.2.3 Punctuation and Segmentation Dataset
@@
3.2.4 Final End-to-End Evaluation Dataset
The final performance of the full cascaded system was assessed on a newly created, representative test set [^1] The final evaluation dataset can be accessed through Hugging Face at https://huggingface.co/datasets/iamTangsang/nepali_to_english_pipeline_evaluation to simulate real-world usage.
@@ 3.2.4 Final End-to-End Evaluation Dataset
  • Questions (60): To test the model’s ability to recognize interrogative tone and punctuation.
@@
3.3 Model Training and Fine-tuning
@@
3.3.1 ASR Model Fine-tuning
@@
The Wav2Vec2 XLS-R utilized the following regularization parameters: attention_dropout  0.1, hidden_dropout  0.1, layerdrop: 0.1. The CTC loss reduction was set to "mean". The key hyperparameters used were as shown in Table 3.
Table 3: Key Hyperparameters for Wav2Vec2.0 Fine-Tuning
@@
Learning Rate (Final) $210^{-5}$
@@
3.3.2 Punctuation Restoration Module Fine-tuning
@@
Table 4: Key Hyperparameters for mT5 Fine-Tuning
@@
3.3.3 NMT Model Multi-Stage Fine-tuning
@@
Table 5: Hyperparameters for MarianMT Pre-training and Fine-tuning
@@
Learning Rate $510^{-5}$ $210^{-5}$
@@ 3.4 Evaluation Metrics
Table 6: Evaluation Types and Corresponding Metrics
@@
3.5 Experimental Setup and Evaluation Scenarios
@@
A dedicated human evaluation was conducted across all three S2TT scenarios using the 900-clip custom test set. Three of the authors conducted a blind human evaluation to assess translation fluency and adequacy on a 1 – 5 Likert scale. The outputs from all system configurations were randomly mixed and anonymized prior to scoring, ensuring that evaluators were unaware of which system configuration produced each translation. After all the ratings were completed, the configuration identities were revealed for analysis. Scores were averaged across raters for each system.
4 Results and Findings

### v2

# arXiv:2602.21647v2[cs.CL] 02 Mar 2026

@@
Abstract.
Cascaded speech-to-text translation (S2TT) systems for low-resource languages can suffer from structural noise, particularly the loss of punctuation during the Automatic Speech Recognition (ASR) phase. This research investigates the impact of such noise on Nepali-to-English translation and proposes an optimized pipeline to mitigate quality degradation. We first establish highly proficient ASR and NMT components: a Wav2Vec2-XLS-R-300m model achieved a state-of-the-art 2.72% CER on OpenSLR-54, and a multi-stage fine-tuned MarianMT model reached a 28.32 BLEU score on the FLORES-200 benchmark. We empirically investigate the influence of punctuation loss, demonstrating that unpunctuated ASR output significantly degrades translation quality, causing a massive 20.7% relative BLEU drop on the FLORES benchmark. To overcome this, we propose and evaluate an intermediate Punctuation Restoration Module (PRM). The final S2TT pipeline was tested across three configurations on a custom dataset. The optimal configuration, which applied the PRM directly to ASR output, achieved a 4.90 BLEU point gain over the direct ASR-to-NMT baseline (BLEU 36.38 vs. 31.48). This improvement was validated by human assessment, which confirmed the optimized pipeline’s superior Adequacy (3.673) and Fluency (3.804) with inter-rater reliability (Krippendorff’s $\alpha\geq 0.723$). This work validates that targeted punctuation restoration is the most effective intervention for mitigating structural noise in the Nepali S2TT pipeline. It establishes an optimized baseline and demonstrates a critical architectural insight for developing cascaded speech translation systems for similar low-resource languages.
Speech-to-Text Translation, Low-Resource Languages, Nepali, ASR, NMT, Nepali-to-English, Punctuation Restoration, Wav2Vec2, MarianMT, Cascaded Pipeline.
journal: TALLIPccs: Computing methodologies Speech recognitionccs: Computing methodologies Machine translationccs: Computing methodologies Language resourcesccs: Computing methodologies Natural language generation
1. Introduction
@@
Nepali, spoken by approximately 19 million native speakers, and an additional 14 million second-language speakers (Eberhard et al., 2024), presents a range of challenges for speech translation. These include limited annotated data, frequent code-switching with English (Gurung, 2019), regional dialect variation, and sociolinguistic features such as complex honorifics without direct English equivalents (e.g., t vs. tpA` both translate as “you” but differ in levels of formality). These characteristics complicate both ASR and MT for Nepali.
End-to-end direct Speech-to-Speech translation (S2ST) models are gaining attention in recent years, for their advantages of architectural simplicity, and potential to reduce information loss, and error propagation (Bentivogli et al., 2021). However, such approaches demand large-scale, parallel speech and text data (Sarim et al., 2025), which makes it impractical for Nepali and most LRLs. In contrast, cascaded pipelines, which first transcribe speech with ASR and then translate it with MT, offer a more practical alternative in low-resource settings. Their modularity allows optimization of each component, resulting in more powerful performance (Sarim et al., 2025). Modern self-supervised ASR models like Wav2Vec2 have shown strong performance even with limited labeled data (Yi et al., 2020), and multilingual NMT models like MarianMT and NLLB can be adapted even with modest amounts of LRL data to achieve good performance (Liu et al., 2020; Verma et al., 2022). Consequently, for low-resource scenarios, cascaded pipelines offer greater data efficiency.
However, cascaded systems introduce their own unique challenges. Error propagation is the most prominent example. A particularly underexplored issue in $\text{Nepali}\leftrightarrow\text{English}$ literature is the lack of punctuation in ASR outputs. While ASR models primarily focus on word accuracy, NMT models are trained on punctuated text, relying on punctuation marks for cues such as sentence boundaries, clause separation, emphasis, disambiguation. Their presence or absence can have significant impact on translation quality (Jwalapuram, 2023).
@@ 2. Related Work
2.1. Automatic Speech Recognition (ASR) for Low-Resource Languages
@@
Traditional ASR approaches for Nepali relied on Hidden Markov Models (HMMs), as in the early work by (Ssarma et al., 2017), which achieved 74.49% accuracy for single-word inputs and 55.55% for three-word phrase inputs Nepali recognition using Voice Activity Detection (VAD), primarily limited by their ability to model complex acoustic variations effectively. The advent of deep learning introduced more sophisticated models like RNN-CTC and CNN-RNN hybrids. (Regmi et al., 2019) implemented an RNN-CTC model trained on a 1,320-word custom dataset, that achieved a 34% CER, but struggled with speaker variability and generalization. (Bhatta et al., 2020) reported 1.83% CER and 11% WER using a CNN-GRU architecture on the OpenSLR43 text-to-speech (TTS) dataset. However, the studio-quality nature of OpenSLR43 limits real-world applicability. Other efforts, such as (Dhakal et al., 2022), utilized bidirectional LSTM paired with ResNet and one-dimensional CNN to achieve a 17.06% CER on OpenSLR dataset. While these studies laid foundational groundwork for Nepali ASR, the models were less capable of generalizing to natural speech variations.
Recent self-supervised models like Wav2Vec2 and Whisper have reshaped ASR for LRLs. By leveraging large-scale unlabeled data for pre-training, they enable better generalization with limited examples (Zhu et al., 2021; Hsu et al., 2024; Fatehi, 2023). Our work builds upon this paradigm by utilizing the Wav2Vec2-XLS-R-300m model on OpenSLR54 and Common Voice v17. This approach enables us to establish a competitive performance benchmark for Nepali ASR.
2.2. Machine Translation (MT) for Low-Resource Languages
Similar to ASR, NMT for LRLs is constrained by lack of high-quality parallel corpora and linguistic complexities. Traditional Statistical Machine Translation (SMT) and rule-based methods dominated early efforts, but transformer-based NMT has since become the standard. However, NMT models require large datasets and are sensitive to rare words, long sentences, domain mismatch and word-alignment issues (Koehn and Knowles, 2017).
To address these limitations, researchers have explored transfer learning, backtranslation, data augmentation, and pivot languages (Haddow et al., 2022; Talwar and Laasri, 2025; Zoph et al., 2016). For instance, MarianMT and NLLB are pretrained multilingual models that support rapid adaptation to LRLs, even with modest amounts of data (Liu et al., 2020; Verma et al., 2022). Recent work by (Verma et al., 2022) demonstrated the effectiveness of multi-stage fine-tuning strategy involving multilingual pre-training followed by language-pair specific fine-tuning, followed by domain fine-tuning.
@@ 2.2. Machine Translation (MT) for Low-Resource Languages
2.3. Challenges in ASR-MT Integration and Punctuation Restoration
Cascaded ASR-MT pipelines, while practical for LRLs, are inherently susceptible to error propagation, where transcription errors negatively affect downstream translation. Since ASR models like Wav2Vec2 primarily focus on word accuracy, and are not inherently trained to predict punctuation, the resultant degradation from directly feeding the unpunctuated ASR output to NMT models is an under-explored issue, especially for LRLs. Punctuation conveys structural and semantic cues like sentence boundaries, emphasis, disambiguation, which NMT models heavily rely on for accurate translation. The work by (Jwalapuram, 2023) quantified the effects of punctuation on MT specifically for German-English, Japanese-English and Ukrainian-English language pairs and concluded that models are heavily sensitive to punctuation.
@@ 3. Methodology
3.1. Overall System Architecture
@@
Figure 1. Cascaded Nepali-to-English Speech-To-Text-Translation System Architecture
3.2. Datasets
@@ 3.2.1. ASR Datasets
Table 1. Nepali ASR datasets used for fine-tuning.
Dataset Total Utterances/Clips Hours (Validated) Used Utterances
Common Voice v17a 742 (validated) + 595 (internally reviewed ‘other’ split) = 1,337 1.0 1,337
OpenSLR54b $\sim$157,000 (initial) $\rightarrow$ 148,188 (post-numerals-filtering) $\sim$143 (pre-5s filter) 136,095 (post-5s filter)
From the Common Voice v17 (ne-NP) dataset, we used 1,337 utterances (approximately 2 hours) from 32 speakers. We combined the ’validated’ split with internally reviewed and validated samples from the ’other’ split for initial fine-tuning. All audio files were uniformly resampled to 16 kHz.
For OpenSLR54, utterances containing numerals were removed to ensure the model outputs numbers in words rather than digits, since inconsistent transcriptions were common. For example, the numeral 007 appeared as ”Ejro Ejro sAt” in one recording, ”sΔ Σy sΔ Σy sAt” in another, and simply ”sAt” elsewhere. To avoid confusing the model, all such utterances were excluded.
@@
3.2.2. NMT Datasets and Multi-Stage Strategy
@@
Table 2. Corpora used at each stage of NMT training and their filtering/rationale.
Stage Corpus Type Size (Pairs) Primary Source / Filtering Rationale
Stage 1: Foundational Filtered Web crawled NLLBa 708,000 Initial 19.6M web-crawled NLLB corpus pairs filtered by LabSE cosine similarity (>0.80>0.80), and removal of Nepali numerals (to maintain consistency). Knowledge Transfer: To learn fundamental translation nuances from filtered, noisy web data.
Stage 2: Expansion Synthetic Corporab,c 5.02M Combination of 1.6M synthetic pairs (Duwal and Bal, 2019) and 3.42M pre-training pairs (Duwal et al., 2024) filtered using a chrF++ cut-off of 50 (translated with 8-bit NLLB/IndicTrans2). Coverage Expansion: Increase exposure to diverse structures and vocabularies using high-volume, quality-filtered synthetic data.
Stage 3: Refinement Manually translated high-Quality corporad,e 210,875 OPUS (GNOME/KDE/Ubuntu), Bible, Global Voices, Penn Treebank + (Acharya and Bal, 2018), Nepal Law Commission, Easy Bible, Nepal Budget Speech 2081/82, Shirish ko Fool Translation. Quality Refinement: Final tuning for translation fluency and accuracy on high-fidelity human-translated pairs.
3.2.3. Punctuation and Segmentation Dataset
@@
3.2.4. Final End-to-End Evaluation Dataset
The final performance of the full cascaded system was assessed on a newly created, representative test set [^1] The final evaluation dataset can be accessed through Hugging Face at https://huggingface.co/datasets/iamTangsang/nepali_to_english_pipeline_evaluation to simulate real-world usage. The final performance of the full cascaded system was assessed on a newly created, representative test set to simulate real-world usage.
@@ 3.2.4. Final End-to-End Evaluation Dataset
  • Questions (60): To test the model’s ability to recognize interrogative punctuation.
@@
3.3. Model Training and Fine-tuning
@@
3.3.1. ASR Model Fine-tuning
@@
The Wav2Vec2 XLS-R utilized the following regularization parameters: attention_dropout  0.1, hidden_dropout  0.1, layerdrop: 0.1. The CTC loss reduction was set to ”mean”. The key hyperparameters used were as shown in Table 3.
Table 3. Key Hyperparameters for Wav2Vec2.0 Fine-Tuning
@@
Learning Rate (Final) $2\times 10^{-5}$
@@
3.3.2. Punctuation Restoration Module Fine-tuning
@@
Table 4. Key Hyperparameters for mT5 Fine-Tuning
@@
3.3.3. NMT Model Multi-Stage Fine-tuning
@@
Table 5. Hyperparameters for MarianMT Pre-training and Fine-tuning
@@
Learning Rate $5\times 10^{-5}$ $2\times 10^{-5}$
@@ 3.4. Evaluation Metrics
Table 6. Evaluation Types and Corresponding Metrics
@@
3.5. Experimental Setup and Evaluation Scenarios
@@
A dedicated human evaluation was conducted across all three S2TT scenarios using the 900-clip custom test set. Three of the authors conducted a blind human evaluation to assess translation fluency and adequacy on a 1 – 5 Likert scale. The outputs from all system configurations were randomly mixed and anonymized prior to scoring, ensuring that evaluators were unaware of which system configuration produced each translation. After all the ratings were completed, the configuration identities were revealed for analysis. Scores were averaged across raters for each system. To validate the consistency of the human assessments, inter-rater reliability was calculated using Krippendorff’s Alpha($\alpha$) for both metric, treating the Likert scores as ordinal data.
4. Results and Findings